OpenStudy (anonymous):

Is it true that Log e^(3x+4) + Log e^(6x+8) is equal to In (3x+4) + In (6x+8) If its true Proved it Please

OpenStudy (shamim):

do u mean e basis

OpenStudy (anonymous):

Yeah .

OpenStudy (shamim):

true

OpenStudy (anonymous):

idk , If that Log e^(f(x)) = In X

OpenStudy (shamim):

no

OpenStudy (shamim):

loge^x=ln x

OpenStudy (anonymous):

That is a derivative Question . y= 1/3 Log e^(3x+4) + Log e^(6x+8) Then I simplified it to y=1/3 In (3x+4) + In (6x+8) is that Correct ?

OpenStudy (shamim):

correct

OpenStudy (anonymous):

Really how come , Can you explain it more ?

OpenStudy (anonymous):

Hows my simplification correct ?

OpenStudy (anonymous):

loge^(3x+4)= you mean this?